OpenStudy (anonymous):

show\[||x||_1\leq \sqrt{n}||x||_2 \text{ in } \mathbb R^n\]

myininaya (myininaya):

\[||x||_1=\sqrt{x_1^2+x_2^2+\cdot \cdot \cdot x_n^2}\] ? what does \[||x||_2=?\]

OpenStudy (anonymous):

\[||x||_2=\sqrt{x_1^2+x_2^2+\cdot \cdot \cdot x_n^2}\] \[||x||_1=\sum|x_k|\]

myininaya (myininaya):

ok so \[||x||_2^2=x_1^2+x_2^2+\cdot \cdot \cdot +x_n^2\] \[||x||_1=\sqrt{x_1^2}+\sqrt{x_2^2}+ \cdot \cdot \cdot \sqrt{x_n^2}\] i'm still thinking...

myininaya (myininaya):

\[||x||_1^2=(\sqrt{x^2_1}+\sqrt{x^2_2}+ \cdot \cdot \cdot +\sqrt{x^n_2})^2\]

OpenStudy (anonymous):

\[||x||_1^2=\sum|x_k|^2+2\sum_{i\neq j}|xi||xj|=||x||_2^2+2\sum_{i\neq j}|xi||xj|\]

myininaya (myininaya):

say the \[\max(x_1,x_2,x_3,...,x_n)=x_i\] so couldn't we say \[||x||^2_2<nx_i^2\] \[||x||^2_1<(n \sqrt{x_i^2})^2\] but \[n x^2 _i < (n \sqrt{x^2_i})^2\] so \[\sqrt{n} x_i< n \sqrt{x^2_i}\] ..... thinking

myininaya (myininaya):

i think you can use the max of the vector in R_n somehow to show this.